Circular No. 8735

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 HR_{30} (SIDING SPRING) Reports from two groups of observers indicate that 2006 HR_{30}, re-

ported as asteroidal when discovered by the Siding Spring Survey (discovery observation tabulated below; cf. MPEC 2006-H40), appears cometary. A co-addition of three 10-s V -band exposures, obtained on July 29.3 UT by S. C. Lowry and A. Fitzsimmons (Department of Physics and Astronomy, Queen’s University, Belfast) with the 3.5-m New Technology Telescope at the European Southern Observatory (ESO), La Silla, show a coma extending to 5′′.5 in p.a. 31^{o}. Observations by M. Hicks and K. Lawrence (Jet Propulsion Laboratory) on Aug. 3 with the Palomar 5-m Hale telescope show a faint coma extending ≈ 10′′ from the ‘nucleus’ in both the R-band and long-slit spectral images.

2006 UT  α_{2000}  δ_{2000}  Mag.
Apr. 20.78181  22^{h}04^{m}31.^{s}90  −9^{o}04′14′′.9  18.7

Additional astrometry, the following orbital elements (from observations in 2005 and 2006), and an ephemeris appear on MPEC 2006-P09. There is also an unconfirmed observation of this object on a 2-hr ESO exposure from 1986 Mar. 5, measured by G. Forti (Arcetri Observatory, Florence) and published on MPS 174727.

Epoch = 2007 Jan. 20.0 TT

T = 2007 Jan. 2.2815 TT  ω = 117.^{o}4137
e = 0.843142  Ω = 309.9512 }2000.0
q = 1.226435 AU  i = 31.8845

a = 7.818745 AU n^{o} = 0.0450815 P = 21.86 years

COMETS C/1999 V5 AND C/1999 X3 (SOHO) Further to IAUC 8734, given below are the discovery data for ad-

ditional small, stellar-appearing comets found on archival SOHO images. C/1995 V5, a Kreutz sungrazer, was of mag ∼ 7.5 or fainter. C/1999 X3, not a member of any known comet group (T = 1999 Dec. 12.7 TT, q = 0.081, e = 1.0, i = 7^{o}.2), was too faint for photometry.

Comet 1999 UT α_{2000} δ_{2000} Inst. F MPEC
C/1999 V5 Nov. 14.379 15^{h}13.^{m}6 −19^{o}34′ C2 HS 2006-O20
C/1999 X3 Dec. 12.564 17 22.9 −23 22 C2 HS 2006-L20
2006 August 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green